### v2

# A multiplicity formula for the Milnor number of smoothable curves

Abstract.
We derive a multiplicity formula for the Milnor number of a reduced smoothable curve singularity generalizing a well-known formula due to Lê, Greuel and Teissier for complete intersection curves. We obtain a multiplicity characterization of Whitney equisingularity for families of locally smoothable curves.
Key words and phrases:
Milnor number, Euler characteristic, delta invariant, complete intersection discrepancy, conormal space, polar varieties, Hilbert–Samuel and Buchsbaum–Rim multiplicities, Whitney equisingularity
2010 Mathematics Subject Classification:
32S15, 32S30, 32S60, 14C17, 13H15
1. Introduction
Let $(X_{0},0)\subset(\mathbb{C}^{n},0)$ be a reduced complex analytic curve singularity. In [BG80] Buchweitz and Greuel define the Milnor number of $(X_{0},0)$ as follows. Let $n\colon\overline{(X_{0},0)}\rightarrow(X_{0},0)$ be the normalization morphism. Denote by $\omega_{X_{0},0}$ the dualizing module of Grothendieck. Denote by $d$ the composition of $\Omega_{X_{0},0}^{1}\rightarrow n_{*}\Omega_{\overline{X_{0},0}}^{1}\cong n_{*}\omega_{\overline{X_{0},0}}\rightarrow\omega_{X_{0},0}$ and the exterior derivation $\mathcal{O}_{X_{0},0}\rightarrow\Omega_{X_{0},0}^{1}$. Then the Milnor number $\mu$ of $(X_{0},0)$ is defined as
$\mu=\mu(X_{0},0):=\dim_{\mathbb{C}}(\omega_{X_{0},0}/d\mathcal{O}_{X_{0},0}).$
In [BG80, Prp. 1.2.1] Buchweitz and Greuel showed that $\mu=2\delta+r-1$, where $\delta:=\dim_{\mathbb{C}}(\overline{\mathcal{O}_{X_{0},0}}/\mathcal{O}_{X_{0},0})$ and $r$ is the number of branches of $(X_{0},0)$. This is a formula discovered for plane curves by Milnor. When $(X_{0},0)$ is smoothable, Bassein [Bas77] proved that $\mu$ equals the first Betti number of the smoothing; in particular, for complete intersection curves $\mu$ coincides with the usual Milnor number.
In [AK70, VIII]) it is shown that $\omega_{X_{0},0}$ is canonically isomorphic to the module of Rosenlicht’s regular differential forms $\omega_{X_{0},0}^{R}$ on $(X_{0},0)$. When the parametrization (normalization) of $(X_{0},0)$ is known, using residues, one can compute $\mu$ in practice. If the parametrization of $(X_{0},0)$ is not known, one would like to have an efficient way for computing $\mu$ (or $\delta$) just from the defining equations of $(X_{0},0)$ in $(\mathbb{C}^{n},0)$.
Denote by $\mathrm{Jac}(X_{0})$ the Jacobian ideal of $(X_{0},0)$ in $\mathcal{O}_{X_{0},0}$, which is the first Fitting ideal of $\Omega_{X_{0},0}^{1}$. Denote by $e(\mathrm{Jac}(X_{0}))$ the Hilbert–Samuel (HS) multiplicity of $\mathrm{Jac}(X_{0})$ and denote by $m$ the multiplicity of $(X_{0},0)$. Because $(X_{0},0)$ is Cohen–Macaulay these multiplicities are easy to compute. Each one of them is the colength of a principal ideal in $\mathcal{O}_{X_{0},0}$ the HS multiplicity of $\mathrm{Jac}(X_{0})$ is the colength of the ideal generated by a general $\mathbb{C}$-linear combination of the $(n-1)\times(n-1)$ minors of the Jacobian matrix of $X_{0}$; the multiplicity of $(X_{0},0)$ is the colength of the ideal generated by a general $\mathbb{C}$-linear combination of generators for the maximal ideal of $\mathcal{O}_{X_{0},0}$. Teissier [T73, Proposition II.1.2] for plane curves and more generally, Lê [Lê74] and Greuel [Gre73] for complete intersection curves, derived the following formula
$\mu=e(\mathrm{Jac}(X_{0}))-m+1.$
We say that $(X_{0},0)$ is smoothable if there exists a flat deformation $h\colon(X,0)\rightarrow(Y,0)$ with $Y$ smooth of dimension one such that $(h^{-1}(0),0)=(X_{0},0)$ and the fiber $h^{-1}(y):=X_{y}$ is smooth for $y\neq 0$ close enough to 00. The main result of this paper generalizes the formula above to the case of smoothable curves by introducing a correction term that loosely speaking measures how far $(X_{0},0)$ is from a complete intersection. Let $(Z_{0},0)\subset(\mathbb{C}^{n},0)$ be a generic complete intersection curve which contains $(X_{0},0)$ as a subvariety. Set $W_{0}:=\overline{Z_{0}\setminus X_{0}}$. Denote by $I_{0}(X_{0},W_{0})$ the intersection number of $X_{0}$ and $W_{0}$ at 00 (see [Hir57, pg. 179]). It is equal to $\dim_{\mathbb{C}}\mathcal{O}_{\mathbb{C}^{n},0}/(I_{X_{0}}+I_{W_{0}})$, where $I_{X_{0}}$ and $I_{W_{0}}$ are the ideals of $X_{0}$ and $W_{0}$ in $\mathcal{O}_{\mathbb{C}^{n},0}$, respectively. We will call $I_{0}(X_{0},W_{0})$ the complete intersection discrepancy of $X_{0}$.
Theorem 1.1.
Suppose $(X_{0},0)\subset(\mathbb{C}^{n},0)$ is a reduced smoothable curve. Then
(1) $\mu=e(\mathrm{Jac}(X_{0}))-I_{0}(X_{0},W_{0})-m+1.$
When $(X_{0},0)$ is a complete intersection, then $I_{W_{0}}=(1)$ and so $I_{0}(X_{0},W_{0})$ vanishes. Thus (1) recovers the Lê–Greuel–Teissier formula.
Let $(X,0)\rightarrow(Y,0)$ be a one-parameter embedded flat deformation of $(X_{0},0)\subset(\mathbb{C}^{n},0)$ with $X_{y}$ smooth for small enough $y\neq 0$. Denote by $Z$ the induced deformation on $Z_{0}$ and define $W$ accordingly. Key to the proof of Theorem 1.1 is interpreting $e(\mathrm{Jac}(X_{0}))$ as the Buchsbaum–Rim multiplicity of a module whose presentation matrix corresponds to the Jacobian matrix of $Z_{0}$. Then Gaffney’s Multiplicity-Polar Theorem implies that $e(\mathrm{Jac}(X_{0}))$ equals to the sum of the covering degree over $(Y,0)$ of the relative polar curve of $(X,0)$ and $I(X_{y},W_{y})$, which by a flatness argument is equal to $I_{0}(X_{0},W_{0})$. By Morse theory, the covering degree of the polar curve is equal to $\chi(X_{y}\cap H)-\chi(X_{y})$, where $H$ is general hyperplane in $\mathbb{C}^{n}$, which is in turn equal to $\mu+m-1$ by the result of Bassein. Combining these equalities we obtain (1).
Whitney equisingularity. Throughout the paper when dealing with complex analytic germs we will work with suitably small representatives of these germs. Let $(X,0)\subset(\mathbb{C}^{n+k},0)$ be a reduced complex analytic space of pure dimension $k+1$, and let $(Y,0)$ be a linear subspace of $(X,0)$ of dimension $k$. Choose an embedding of $(X,0)$ in $\mathbb{C}^{n+k}=\mathbb{C}^{n}\times\mathbb{C}^{k}$, so that $(Y,0)$ is represented by $0\times U$, where $U$ is an open neighborhood of 00 in $\mathbb{C}^{k}$. Let $\mathrm{pr}:\mathbb{C}^{n}\times\mathbb{C}^{k}\rightarrow\mathbb{C}^{k}$ be the projection on the second factor and let $i:\mathbb{C}^{k}\rightarrow\mathbb{C}^{n}\times\mathbb{C}^{k}$ be the inclusion $i(y)=(0,y)$. Denote by $\pi_{Y}$ be the retraction $i\circ\mathrm{pr}_{|X}$. View $X$ as the total space of the family $\pi_{Y}:(X,0)\rightarrow(Y,0)$. For each $y\in Y$ set $X_{y}:=X\cap\pi^{-1}_{Y}(y)$. We say $H$ is a tangent hyperplane at $x\in X-Y$ if $H$ is a hyperplane in $\mathbb{C}^{n+k}$ that contains the tangent space $T_{x}X$. Let $(x_{i})$ be a sequence of points from $X-Y$ and $(y_{i})$ be a sequence of points from $Y$ both converging to 00. Suppose that the sequence of secants $(\overline{x_{i}y_{i}})$ has limit $l$ and the sequence of tangent hyperplanes $\{T_{x_{i}}X\}$ has limit $T$. We say that $(X,0)\rightarrow(Y,0)$ is Whitney equisingular at 00, or that the pair of strata $(X-Y,Y)$ satisfies the Whitney conditions at 00 if $l\subset T$.
By a result of Lê and Teissier [LT88] $(X,0)\rightarrow(Y,0)$ is Whitney equisingular if and only if $X_{y}$ is equimultiple along $Y$ and there exists a homeomorphism $g\colon(\mathbb{C}^{n+k},0)\rightarrow(\mathbb{C}^{n+k},0)$ whose restriction to $X$ induces the homeomorphism $(X,0)\cong(X_{0}\times Y,0)$. For families of reduced curves it is shown in [BGG80, Thm. III.3] that $(X,0)\rightarrow(Y,0)$ is Whitney equisingular if and only if the Milnor number $\mu(X_{y})$ and the multiplicity $m(X_{y})$ at 00 are constant along $Y$.
We say that $X_{y}$ is locally smoothable if the germ of $X_{y}$ at each of its singular points is smoothable. Denote by $e(\mathrm{Jac}(X_{y},0))$ the HS-multiplicity of the Jacobian ideal of the germ $(X_{y},0)$. As a direct consequence of Theorem 1.1 and [BGG80, Thm. III.3] we have the following corollary.
Corollary 1.2.
Assume that the fibers $X_{y}$ are reduced curves with locally smoothable singularities. The following holds:
  1. (i)
    Assume $e(\mathrm{Jac}(X_{y},0))-I_{0}(X_{y},W_{y})$ is independent of $y$ near 00. Then the union of the singular points of $X_{y}$ is $Y$ and the pair $(X-Y,Y)$ satisfies the Whitney conditions at 00.
  2. (ii)
    Suppose $(X-Y,Y)$ satisfies the Whitney conditions at 00. Then $e(\mathrm{Jac}(X_{y},0))-I_{0}(X_{y},W_{y})$ is independent of $y$ near 00.
Acknowledgements. The authors would like to thank Steven Kleiman and Bernard Teissier for helpful and stimulating discussions. The first author was supported by the Bulgarian Ministry of Education and Science Scientific Programme ”Enhancing the Research Capacity in Mathematical Sciences (PIKOM)”, No. DO1-67/05.05.2022. The third author was supported by a “Petar Beron i NIE” fellowship [KP-06-D15-1] from the Bulgarian Science Fund and was partially funded by a Marie Składowska-Curie fellowship GTSP-$101111114$ from the European Commission.
2. Conormal spaces, polar varieties, and multiplicities
2.1. Conormal spaces
Let $h\colon(X,0)\rightarrow(Y,0)$ be a flat embedded deformation of a reduced complex analytic curve $(X_{0},0)\subset\mathbb{C}^{n}$ with $(Y,0)$ smooth of dimension one. First we describe the relative conormal variety $C_{\mathrm{rel}}(X)$ of $X$ in $\mathbb{C}^{n+1}$ using the relative Jacobian module of $X$. Suppose $X$ is reduced. Let $X$ be defined by the vanishing of some analytic functions $f_{1},\ldots,f_{p}$ on a Euclidean neighborhood of 00 in $\mathbb{C}^{n+1}$. Consider the following conormal exact sequence
[…]
where $I$ is the ideal of $X$ in $\mathcal{O}_{\mathbb{C}^{n}\times Y,0}$ and the map $\delta$ sends a function $f$ vanishing on $X$ to its differential $df$. Dualizing we obtain the following nested sequence of torsion-free sheaves:
$\mathrm{Image}(\delta^{*})\subset(\mathrm{Image}\ \delta)^{*}\subset(I/I^{2})^{*}.$
Observe that locally the sheaf $\mathrm{Image}(\delta^{*})$ can be viewed as the column space of the relative Jacobian matrix of $X$. If $(x_{1},\ldots,x_{n},t)$ are coordinates on $(\mathbb{C}^{n+1},0)$, where $t$ is the $Y$-coordinate, then the relative Jacobian matrix is simply the $p$ by $n$ matrix $(\partial f_{i}({\bf x},t)/\partial x_{j})$.
Denote $\mathrm{Image}(\delta^{*})$ by $J_{\mathrm{rel}}(X)$ and call it the relative Jacobian module of $X$. It’s a module contained in the free module $\mathcal{O}_{X,0}^{p}$. Define the Rees algebra $\mathcal{R}(J_{\mathrm{rel}}(X))$ to be the subalgebra of $\mathrm{Sym}(\mathcal{O}_{X,0}^{p})$ generated by $J_{\mathrm{rel}}(X)$ placed in degree $1$. Define the relative conormal space $C_{\mathrm{rel}}(X)$ of $X$ to be the closure in $X\times\check{\mathbb{P}}^{n-1}$ of the set of pairs $(x,H)$ where $x$ is a smooth point of $X_{h^{-1}(h(x))}$ and $H$ is a (tangent) hyperplane in $\mathbb{C}^{n}$ at $x$ containing $T_{x}X_{h^{-1}(h(x))}$. We have
$C_{\mathrm{rel}}(X)=\mathrm{Projan}(\mathcal{R}(J_{\mathrm{rel}}(X))).$
Indeed, observe that both sides are equal over the smooth part of the fibers $(X_{y})_{\mathrm{sm}}$ to the set of pairs $(x,H)$ where $H$ is a tangent hyperplane in $\mathbb{C}^{n}$ to the simple point $x\in(X_{y})_{\mathrm{sm}}$. The left side is the closure of this set, and so is the right side simply because the Rees algebra is by construction a subalgbebra of the symmetric algebra $\mathrm{Sym}(\mathcal{O}_{X,0}^{p})$.
2.2. Polar varieties
Consider the following diagram
[…]
The dimension of the generic fiber of $\pi$ is $n-2$, so $\dim C_{\mathrm{rel}}(X)$ is $n$. By Kleiman’s transversality theorem [Kl74] (take the transitive action $PGL(n,\mathbb{C})$ on ˇ$\check{\mathbb{P}}^{n-1}$ considered as the set of hyperplanes of $\mathbb{C}^{n}$), for a generic point $H\in\check{\mathbb{P}}^{n-1}$, the pullback $\lambda^{-1}(H)$ is going to be of dimension at most $\dim C_{\mathrm{rel}}(X)-\text{codim}(H)=n-(n-1)=1$. The projection to $X$ of this curve $\Gamma_{\mathrm{rel}}^{1}(X):=\pi(\lambda^{-1}(H))$ is called the relative polar variety of dimension one, or simply the relative polar curve of $h\colon(X,0)\rightarrow(Y,0)$. Again by Thm. 2 (ii) and Rmk. 7 in [Kl74], choosing $H$ generic enough guarantees that $\Gamma_{\mathrm{rel}}^{1}(X)$ is reduced because by assumption $X$ is reduced. An equivalent way to obtain $\Gamma_{\mathrm{rel}}^{1}(X)$ is to consider a linear form $F$ on $\mathbb{C}^{n}$ that defines a hyperplane that is not a limiting tangent hyperplane to $X_{0}$ at 00. Then $\Gamma_{\mathrm{rel}}^{1}(X)$ is the closure of the set of critical points of $F$ restricted to the smooth part of the fibers of $h$ ([FT18, Prp. 3.4]).
Suppose $h\colon(X,0)\rightarrow(Y,0)$ is an embedded deformation of $(X_{0},0)\subset(\mathbb{C}^{n},0)$, which is a smoothing of $(X_{0},0)$. Let $F\colon(\mathbb{C}^{n},0)\rightarrow(\mathbb{C},0)$ be a general linear functional. Denote by $\mathrm{deg}_{Y}\Gamma_{\mathrm{rel}}^{1}(X)$ the number of points in the fiber $\Gamma_{\mathrm{rel}}^{1}(X)_{y}$ for $y\neq 0$ generic. Denote by $m$ the multiplicity of $(X_{0},0)$.
Proposition 2.1.
We have the following Euler characteristic-degree relation.
(3) $\mathrm{deg}_{Y}\Gamma_{\mathrm{rel}}^{1}(X)=\chi(X_{y}\cap F^{-1}(0))-\chi(X_{y})=\mu+m-1.$
Proof.
Assume $(X,0)\subset(\mathbb{C}^{n+1},0)$. Let $X_{0}$ be embedded in a small open ball $B_{0}:=B(0,\epsilon)\subset(\mathbb{C}^{n},0)$. Replace $Y$ by a small enough open disk centered at 00. We may assume that $X$ is a closed analytic subset of $B_{0}\times Y$ and we can assume that $X$ is smooth outside the origin. Let $(x_{1},\ldots,x_{n},t)$ be coordinates for $(\mathbb{C}^{n+1},0)$. Let $F\colon(\mathbb{C}^{n+1},0)\rightarrow(\mathbb{C},0)$ be the linear functional $(x_{1},\ldots,x_{n},t)\rightarrow(\sum_{i=1}^{n}\alpha_{i}x_{i})$ where $\alpha_{i}$ are complex numbers subject to finitely many genericity conditions specified below. Denote by $\Sigma_{X}(F)$ and by $\Sigma_{X_{y}}(F)$ the critical loci of $F$ on $X$ and $X_{y}$, respectively. Note that by ([FT18, Prp. 3.4]) $\Gamma_{\mathrm{rel}}^{1}(X)$ is the closure of $\Sigma_{X_{y}}(F)$ for $y\neq 0$ close enough to 00. Thus $\mathrm{deg}_{Y}\Gamma_{\mathrm{rel}}^{1}(X)=\#(\Sigma_{X_{y}}(F))$. Because $\Gamma_{\mathrm{rel}}^{1}(X)$ meets $\partial B_{0}\times Y$ at finitely many points, for $y$ close enough to 00, the functional $F_{|X_{y}}$ does not have critical points on $S_{\epsilon}^{2n-1}$. Note that in case $Y=0\times\mathbb{C}\subset X$, then by choosing the $\alpha_{i}$s generic enough we can guarantee that $\Sigma_{X}(F)\cap(Y,0)=0$ and thus $0\in\mathbb{C}^{n}$ is not in $\Sigma_{X_{y}}(F)$ for small enough $y\neq 0$. By [NOT13, Thm. A.5] (cf. [Kav04, Thm. 1.4])
#$\#(\Sigma_{X_{y}}(F))=\chi(F^{-1}(0)\cap X_{y})-\chi(X_{y}).$
By [Bas77] (cf. [BG80, Cor. 4.2.3]) $\mu(X_{0},0)=1-\chi(X_{y})$. Thus it remains to show that $\chi(F^{-1}(0)\cap X_{y})=m$.
Set $f:=\sum_{i=1}^{n}\alpha_{i}x_{i}$. For each $y\in Y$ denote by $f_{y}$ the image of $f$ in $\mathcal{O}_{X_{y}}$. Because $(X_{0},0)$ is a Cohen–Macaulay curve, for generic $\alpha_{i}$ we have $m=\dim_{\mathbb{C}}\mathcal{O}_{X_{0},0}/f_{0}$. Because $h$ is flat and $X_{0}$ is reduced, then $(t,f)$ is a regular sequence in $\mathcal{O}_{X,0}$, and so is $(f,t)$, because $\mathcal{O}_{X,0}$ is local. By taking sufficiently small representatives of the germs $(X,0)$ and $(Y,0)$ we can ensure that $\mathcal{O}_{X,0}/f$ is a finite $\mathcal{O}_{Y,0}$-module (see [GrR84, Prp. 1), p. 63]). Because $\mathcal{O}_{X,0}/f$ is a torsion-free $\mathcal{O}_{Y,0}$-module, we have $\dim_{\mathbb{C}}\mathcal{O}_{X_{0},0}/f_{0}=\dim_{\mathbb{C}}\mathcal{O}_{X_{y}}/f_{y}.$ But $\dim_{\mathbb{C}}\mathcal{O}_{X_{y}}/f_{y}=\#(F^{-1}(0)\cap X_{y})=\chi(F^{-1}(0)\cap X_{y})$. Thus $\chi(F^{-1}(0)\cap X_{y})=m$. ∎
2.3. Multiplicities
Let $(V,0)$ be a reduced complex analytic variety. Let $M$ be an $\mathcal{O}_{V}$-module such that its rank at the generic point of each irreducible component of $V$ is $e$. Assume $M$ is contained in a free module $\mathcal{O}_{V}^{p}$ for some $p$. Denote by $[M]\in\mathrm{Mat}(p\times c,\mathcal{O}_{V})$ a presentation matrix of $\mathcal{O}_{V}^{p}/M$. Let $A\in\text{Mat}(e\times p,\mathbb{C})$ be a generic matrix such that $[M_{e}]:=A[M]$ has rank $e$ at the generic point of each irreducible component of $(V,0)$. Let $M_{e}$ be the $\mathcal{O}_{V}$-module generated by the columns of $[M_{e}]$. Then $M_{e}$ is contained in a free module $F:=\mathcal{O}_{V}^{e}$. Denote by $\mathcal{R}(M)$ and $\mathcal{R}(M)$ the Rees algebras of $M$ and $M_{e}$, respectively.
Proposition 2.2.
We have
$\mathrm{Projan}(\mathcal{R}(M))=\mathrm{Projan}(\mathcal{R}(M_{e})).$
Proof.
By construction the row spaces of $[M]$ and of $[M_{e}]$ agree on the Zariski open dense subset of $V$ where the rank of these matrices is $e$. Over this Zariski open set the spaces $\mathrm{Projan}(\mathcal{R}(M))$ and $\mathrm{Projan}(\mathcal{R}(M_{e}))$ are just the projectivization of these row spaces, hence they are equal. Since the spaces agree on a Zariski open dense set they agree everywhere. ∎
Suppose now that $(V,0)$ is a reduced curve. Then $\mathrm{Supp}_{V}(F/M_{e})=\{0\}$. The inclusion $M_{e}\subset F$ induces an inclusion of the Rees algebra $\mathcal{R}(M_{e})$ in the symmetric algebra $\mathrm{Sym}(F)$. Denote by $M_{e}^{l}$ and $F^{l}$ the $l$th graded components of $\mathcal{R}(M_{e})$ and $\mathrm{Sym}(F)$, respectively. Then by results of [BR64] for $l$ large enough
$\dim_{\mathbb{C}}F^{l}/M_{e}^{l}=e(M_{e},F)l^{e}/e!+O(l^{e-1}).$
The coefficient $e(M_{e},F)$ is called the Buchsbaum–Rim (BR) multiplicity of the pair $(M_{e},F)$. The BR-multiplicity generalizes the HS-multiplicity of ideals to modules. A submodule $M_{e}^{\prime}$ of $M_{e}$ is called a reduction of $M_{e}$ if $\mathcal{R}(M_{e})$ is an integral extension of $\mathcal{R}(M_{e}^{\prime})$. By [SH06, Cor. 16.4.7] there exists a reduction of $M_{e}$ generated by $e$ generic linear combinations of the columns of $[M_{e}]$, or in other words for a generic matrix $B\in\text{Mat}(c\times e,\mathbb{C})$ such that the module generated by the columns of $[M_{e}]B$ is a reduction of $M_{e}$. By [SH06, Cor. 16.5.7] $e(M_{e},F)=e(M_{e}^{\prime},F)$. Because $(V,0)$ is Cohen–Macaulay by [BR64, 4.5, p. 223] $e(M_{e}^{\prime},F)=\dim_{\mathbb{C}}(\mathcal{O}_{V,0}/\mathrm{Fitt}_{0}(F/M_{e}^{\prime}))$, where $\mathrm{Fitt}_{0}(F/M_{e}^{\prime})$ is the determinant of the $e\times e$ matrix $[M_{e}^{\prime}]$.
The Fitting ideal $\mathrm{Fitt}_{p-e}(\mathcal{O}_{V}^{p}/M)$ is generated by the $e\times e$ minors of $[M]$. It’s an ideal in $\mathcal{O}_{V,0}$ primary to the maximal ideal; thus it has a well-defined HS-multiplicity $e(\mathrm{Fitt}_{p-e}(\mathcal{O}_{V}^{p}/M))$.
Proposition 2.3.
For a generic choice of $A$ we have $e(M_{e},F)=e(\mathrm{Fitt}_{p-e}(\mathcal{O}_{V}^{p}/M)).$
Proof.
Recall that $[M]\in\mathrm{Mat}(p\times c,\mathcal{O}_{V})$. By $K$ and $L$ we will denote $e$-element subsets of $\{1,\ldots,p\}$ and $\{1,\ldots,c\}$, respectively. Denote by $[M]_{K,L}$ the $e\ \times e$ minor of $[M]$ with $K$ rows and $L$ columns. Denote by $A_{K}$ the $e\times e$ minor of $A$ with $K$ columns, and denote by $B_{L}$ the $e\times e$ minor with $L$ rows. By construction $[M_{e}^{\prime}]=A[M]B$. Applying the Cauchy–Binet formula we obtain
$\mathrm{det}[M_{e}^{\prime}]=\sum_{K,L}A_{K}[M]_{K,L}B_{L}.$
Note that $[M]_{K,L}$ give the $e\times e$ minors of $[M]$. Set $I:=\mathrm{Fitt}_{p-e}(\mathcal{O}_{V}^{p}/M)$. The exceptional divisor $D$ of $\mathrm{Bl}_{I}V$ is set-theoretically a finite set of points in $\mathbb{P}^{l-1}$. Let $[\alpha_{K,L}]\in\mathbb{P}^{l-1}$ be such a point. In order for a general linear combination $\sum_{K,L}\beta_{K,L}M_{K,L}$ to be a reduction, one needs to have $\alpha\cdot\beta:=\sum_{K,L}\alpha_{K,L}\beta_{K,L}\neq 0$ for all $[\alpha_{K,L}]\in D$. The linear combinations that come from matrices $A$ and $B$ are of the form $\beta_{K,L}=A_{K}\cdot B_{L}$ so in fact we have the following generic condition on $A$ and $B$
(4) $\sum_{K,L}\alpha_{K,L}A_{K}B_{L}\neq 0.$
The sum is a polynomial expression $P(A,B)$ in the entries of $A$ and $B$ so it is enough to show that $P(A,B)\not\equiv 0$, or that there exists a pair of matrices $(A_{0},B_{0})$ such that $P(A_{0},B_{0})\neq 0$. Without loss of generality suppose that $\alpha_{K_{0},L_{0}}\neq 0$ for $K_{0}=L_{0}=\{1,\ldots,e\}$. Set $A_{0}=(\delta_{i,j})_{1\leq i\leq e,1\leq j\leq p}$ and $B_{0}=(\delta_{s,t})_{1\leq s\leq g,1\leq t\leq e}$, where $\delta_{p,q}$ is the Kronecker delta symbol. A direct computation of the minors gives
$(A_{0})_{K}=\left\{\begin{array}[]{cc}1&\text{if }K=K_{0}\\ 0&\text{otherwise}\end{array}\right.\quad\mathrm{and}\quad(B_{0})_{L}=\left\{\begin{array}[]{cc}1&\text{if }L=L_{0}\\ 0&\text{otherwise}.\end{array}\right.$
Thus $P(A_{0},B_{0})=\alpha_{K_{0},L_{0}}\neq 0$ and so (4) holds for generic $A$ and $B$. Intersecting finitely many open sets in $\text{Mat}(e\times p,\mathbb{C})$ and $\text{Mat}(c\times e,\mathbb{C})$ we get that (4) holds for each $[\alpha_{K,L}]\in D$ and for generic $A$ and $B$. Therefore, for such $A$ and $B$ the ideal generated by $\mathrm{det}[M_{e}^{\prime}]$ is a reduction of $\mathrm{Fitt}_{p-e}(\mathcal{O}_{V}^{p}/M)$. By [SH06, Prp. 11.1.10] $e(\mathrm{Fitt}_{p-e}(\mathcal{O}_{V}^{p}/M))=\dim_{\mathbb{C}}\mathcal{O}_{V,0}/(\mathrm{det}[M_{e}^{\prime}]).$ But as observed above
$e(M_{e},F)=\dim_{\mathbb{C}}(\mathcal{O}_{V,0}/\mathrm{Fitt}_{0}(F/M_{e}^{\prime}))=\dim_{\mathbb{C}}\mathcal{O}_{V,0}/(\mathrm{det}[M_{e}^{\prime}]).$
Therefore, $e(M_{e},F)=e(\mathrm{Fitt}_{p-e}(\mathcal{O}_{V}^{p}/M)).$
To get the equality in Proposition 2.3 selecting $A$ so that Proposition 2.2 holds is not enough as indicated in the $(3,4,5)$ example in Section 5. Finally, we remark that $e(M_{e},F)$ was considered in a more general setting in [KT94, Sct. 7]. To define $e(M_{e},F)$ from the point of view of Kleiman and Thorup, one intersects $\mathrm{Projan}(\mathrm{Sym}(\mathcal{O}_{V}^{p}))=V\times\mathbb{P}^{p-1}$ with $p-e$ general hyperplanes from $\mathbb{P}^{p-1}$, blowing up this intersection with the image of the ideal generated by $M$ in $\mathrm{Sym}(\mathcal{O}_{V}^{p})$, and then deriving $e(M_{e},F)$ as a sum of certain intersection numbers of the exceptional divisor.
3. A flatness result
Let $(X_{0},0)\subset(\mathbb{C}^{n},0)$ be a reduced complex analytic curve which is not a complete intersection. Let $(Z_{0},0)\subset(\mathbb{C}^{n},0)$ be a complete intersection which contains $(X_{0},0)$ as a subvariety. Let $(X,0)\rightarrow(Y,0)$ be a flat embedded deformation of $(X_{0},0)$ with $(Y,0)$ smooth of dimension one. Denote by $Z$ the induced deformation on $(Z_{0},0)$ from $(X,0)\rightarrow(Y,0)$. Set $W:=\overline{Z\setminus X}$ and $S:=X\times_{\mathbb{C}^{n}\times Y}W$. In this section we prove that $S$ is flat over $Y$, and in particular, $I(X_{y},W_{y}):=\sum_{s_{y}\in S_{y}}I_{s_{y}}(X_{y},W_{y})$ is constant along $Y$. In Section 4 we will show that if $(X,0)\rightarrow(Y,0)$ is a smoothing of $(X_{0},0)$, then $S$ is also reduced for a generic choice of $Z_{0}$. We begin with an algebraic result.
Lemma 3.1.
Assume $(R,\mathfrak{m})$ is a local Noetherian ring with two ideals $I$ and $J$ such that $\mathrm{depth}(R/I\cap J)\geqslant 2,\,\mathrm{depth}(R/I)\geqslant 1$ and $\mathrm{depth}(R/J)\geqslant 1$. Then $\mathfrak{m}\notin\mathrm{Ass}_{R}(R/I+J)$.
Proof.
Phrased in terms of local cohomology the conclusion of our statement is equivalent to $H_{\mathfrak{m}}^{0}(R/I+J)=0$. Consider the short exact sequence
[…]
Extract from the long exact sequence of local cohomology the following exact sequence
[…]
By the vanishing theorem of local cohomology (see for instance [BS12, Cor. 6.2.8]), we have
$H_{\mathfrak{m}}^{0}(R/I\oplus R/J)=0\quad\text{and}\quad H_{\mathfrak{m}}^{1}(R/I\cap J)=0.$
Therefore, $H_{\mathfrak{m}}^{0}(R/I+J)=0$. ∎
Proposition 3.2.
In the setup above, $S$ is flat over $Y$. In particular, $I(X_{y},W_{y})$ is constant for all $y$ close to 00.
Proof.
We apply Lemma 3.1 with $R=\mathcal{O}_{\mathbb{C}^{n}\times Y,0}$, $I:=I_{X}$, $J:=I_{W}$, and $\mathfrak{m}$ the ideals of $X$, $W$ and 00 in $\mathcal{O}_{\mathbb{C}^{n}\times Y,0}$, respectively. Note that $I_{X}\cap I_{W}=I_{Z}$ and the ideal of $S$ in $\mathcal{O}_{\mathbb{C}^{n}\times Y,0}$ is $I_{X}+I_{W}$. Let $t$ be the uniformizing parameter of $(\mathcal{O}_{Y},0)$. Observe that the depth hypothesis of Lemma 3.1 are satisfied. Indeed, $Z$ is a complete intersection of dimension $2$, so $\mathrm{depth}(\mathcal{O}_{Z})=2$; $X$ is flat over $Y$, and so $t$ is a nonzero divisor of $\mathcal{O}_{X}$ and thus $\mathrm{depth}(\mathcal{O}_{X})\geq 1$. Also, because $I_{W}$ is obtained from the primary decomposition of $I_{Z}$ by removing some of its minimal primes, then $t$ avoids the minimal primes of $I_{W}$, and so $W$ is flat over $Y$. Thus, $\mathrm{depth}(\mathcal{O}_{W})\geq 1$. If $S$ is empty there is nothing to prove. So we can assume that either $S$ is set-theoretically equal to 00, or $S$ is of dimension one. The first case is impossible because by Lemma 3.1 $\mathfrak{m}\notin\mathrm{Ass}(\mathcal{O}_{S})$.
Suppose $S$ has a vertical component, i.e. $X_{0}$ and $W_{0}$ share an irreducible component. Identify $I_{X}$ and $I_{W}$ with their images in $\mathcal{O}_{Z}$. Suppose that as sets $X_{0}\subset W_{0}$. But $Z_{0}=X_{0}\cup W_{0}$, so $W_{0}=Z_{0}$. This means that for each $f\in I_{W}$ we have $f=tf_{1}$ with $f_{1}\in\mathcal{O}_{Z}$. Because $\mathcal{O}_{W}$ is flat over $\mathcal{O}_{Y}$, we have $f_{1}\in I_{W}$. So $I_{W}=tI_{W}$ in $\mathcal{O}_{Z}$. By Nakayama’s lemma $W=Z$, which is impossible.
Assume that $X_{0}$ is not contained in $W_{0}$ set-theoretically. Then there exists an irreducible component of $X_{0}$, which is not a component of $W_{0}$. Consider the union of all such components of $X_{0}$ that are not components of $W_{0}$ and denote it by $X_{0}^{\prime}$. Pick an element $r\in\mathcal{O}_{Z}$ such that its image $\overline{r}$ in $\mathcal{O}_{Z_{0}}:=\mathcal{O}_{Z}/t\mathcal{O}_{Z}$ is in the ideal of $X_{0}^{\prime}$ but not in the ideal of $X_{0}$. Then for each $f\in I_{W}$, there exists $f_{1}\in\mathcal{O}_{Z}$ such that $rf=tf_{1}$. Again, because $\mathcal{O}_{W}$ is flat over $\mathcal{O}_{Y}$, we have $f_{1}\in I_{W}$. Thus $rI_{W}\subset(t)I_{W}$. By the Cayley–Hamilton theorem $\det(A)I_{W}=0$ where $A=(\delta_{i,j}r-a_{i,j})$ is a square matrix of size $m\times m$ for some $m$, $\delta_{i,j}$ is Kronecker’s delta function and $a_{i,j}\in(t)$. But $I_{X}=((0)\colon_{\mathcal{O}_{Z}}I_{W})$. Thus $\det(A)\in I_{X}$. Denote by $\tilde{r}$ the image of $r$ in $\mathcal{O}_{X_{0}}$. Because $\det(A)$ is a monic polynomial in $r$ with coefficients in $(t)$, we have $\tilde{r}^{m}=0$ in $\mathcal{O}_{X_{0}}$. But this is impossible because we chose $r$ so that it’s not in the nilradical of $\mathcal{O}_{X_{0}}$. We reached a contradiction with the assumption that $S$ has a vertical component. Therefore, by taking sufficiently small representatives of the germs $(X,0)$ and $(Y,0)$ we can ensure that $(S,0)\rightarrow(Y,0)$ is finite (see [GrR84, Prp. 1), p. 63]).
Again, by Lemma 3.1, the origin cannot be an embedded component of $S$. Thus, $t$ is not a zero divisor of $\mathcal{O}_{S}$. So, $(S,0)\rightarrow(Y,0)$ is flat, and thus $y\rightarrow I(X_{y},W_{y})$ is constant for all $y$ close to 00. ∎
4. Proofs
Let $h\colon(X,0)\rightarrow(Y,0)$ be a one-parameter embedded flat deformation of $(X_{0},0)\subset(\mathbb{C}^{n},0)$ with $X_{y}$ smooth for small enough $y\neq 0$. Let $(Z_{0},0)\subset(\mathbb{C}^{n},0)$ be a complete intersection curve which contains $(X_{0},0)$ as a subvariety. Denote by $Z$ the induced deformation on $Z_{0}$. If $(X,0)$ is cut out from $(\mathbb{C}^{n}\times Y,0)$ by the vanishing of the analytic functions $f_{1},\ldots,f_{p}$, then the equations for $Z$ are given by $A(f_{1},\ldots,f_{p})^{T}$ for some $A\in\mathrm{Mat}((n-1)\times p,\mathbb{C})$. Set $W:=\overline{Z\setminus X}$ and set $S:=X\times_{\mathbb{C}^{n}\times Y}W$. Observe that because $S$ is flat over $Y$ by Prp. 3.2, we have $W_{y}=\overline{Z_{y}\setminus X_{y}}$ for each $y\in Y$, i.e. the formation of $W$ specializes with passage to the fibers.
Denote by $J_{\mathrm{rel}}(X)$ the relative Jacobian module and denote by $J_{\mathrm{rel}}(Z):=AJ_{\mathrm{rel}}(X)$ the $\mathcal{O}_{X}$-module generated by the relative Jacobian matrix of $Z$. For each $y$ denote by $J(Z_{y})$ the image of $J_{\mathrm{rel}}(Z)$ in $\mathcal{O}_{X_{y}}^{n-1}$. It’s the $\mathcal{O}_{X_{y}}$-module generated by the columns of the Jacobian matrix of $Z_{y}$. We will repeatedly use the fact that the formation of Fitting ideals commutes with base change. In particular,
$\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{n-1}/J_{\mathrm{rel}}(Z))\otimes\mathcal{O}_{X_{y}}=\mathrm{Fitt}_{0}(\mathcal{O}_{X_{y}}^{n-1}/J(Z_{y})).$
Proposition 4.1.
Let $X_{y}$ be a smooth fiber and let $s_{y}\in S_{y}$. Let $u$ be a uniformizing parameter of $\mathcal{O}_{X_{y},s_{y}}$. Suppose
(5) $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{n-1}/J_{\mathrm{rel}}(Z))\otimes\mathcal{O}_{X_{y},s_{y}}=(u).$
Then $(Z_{y},s_{y})$ is a normal crossing divisor in $(\mathbb{C}^{2},s_{y})$. In particular, $I_{s_{y}}(X_{y},W_{y})=1$.
Proof.
[…]
$\mathrm{Fitt}_{0}(\mathcal{O}_{X_{y},s_{y}}^{n-1}/J(Z_{y}))=(\partial{xg(u,x)}/\partial{u},\partial{xg(u,x)}/\partial{x})_{|x=0}=(g(u,0)).$
By assumption $\mathrm{Fitt}_{0}(\mathcal{O}_{X_{y},s_{y}}^{n-1}/J(Z_{y}))=(u)$, thus $\mathrm{ord}_{u}(g(u,0))=1$. This implies that
$\deg(\mathrm{in}(g(u,x)))=1$ and so $(W_{y},s_{y})$ is smooth at $s_{y}$. Also $\dim_{\mathbb{C}}\mathcal{O}_{\mathbb{C}^{2},s_{y}}/(x,g(x,u))=1$. So $(X_{y},s_{y})$ and $(W_{y},s_{y})$ intersect transversally at $s_{y}$. ∎
Let $M\subset\mathcal{O}_{X}^{p}$ be a finitely generated $\mathcal{O}_{X}$-module which is locally free of rank $e$ at the generic point of each irreducible component of $X$. For $A\in\text{Mat}(e\times p,\mathbb{C})$ set $[M_{e}]:=A[M]$, where $[M]$ is a presentation matrix of $\mathcal{O}_{X}^{p}/M$. Let $M_{e}\subset\mathcal{O}_{X}^{e}$ be the $\mathcal{O}_{X}$-module generated by the columns of $[M_{e}]$. Note that by the Cauchy-Binet formula $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{e}/M_{e})\subset\mathrm{Fitt}_{p-e}(\mathcal{O}_{X}^{p}/M)$. Let $S^{\prime}$ be the analytic subspace of $X$ defined by $(\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{e}/M_{e}):_{\mathcal{O}_{X}}\mathrm{Fitt}_{p-e}(\mathcal{O}_{X}^{p}/M))$.
Proposition 4.2.
Suppose $A$ is generic. Then $S^{\prime}$ is generically reduced.
Proof.
Let $\eta$ be the generic point of an irreducible component of $X$. Because $X$ is reduced, $\mathcal{O}_{X,\eta}$ is a field and so $M_{\eta}:=M\otimes\mathcal{O}_{X,\eta}$ and $(M_{e})_{\eta}:=M_{e}\otimes\mathcal{O}_{X,\eta}$ are $\mathcal{O}_{X,\eta}$-vector spaces. So $\mathcal{O}_{X,\eta}^{p}/M_{\eta}$ is $\mathcal{O}_{X,\eta}$-vector space of dimension $p-e$ and $\mathcal{O}_{X,\eta}^{e}=(M_{e})_{\eta}$. Thus the closed subsets of $X$ defined by $\mathrm{Fitt}_{p-e}(\mathcal{O}_{X}^{p}/M)$ and $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{e}/M_{e})$ are nowhere dense in $X$. Therefore, $S^{\prime}$ is either empty or of dimension $1$.
Suppose $[M]\in\text{Mat}(p\times l,\mathcal{O}_{X})$. Denote by $M^{*}$ be the $\mathcal{O}_{X}$-submodule of $\mathcal{O}_{X}^{l}$ generated by the columns of $[M]^{\mathrm{tr}}$. We have $\mathrm{Fitt}_{p-e}(\mathcal{O}_{X}^{p}/M)=\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{*})$. Denote by $M^{\prime}$ the $\mathcal{O}_{X}$-submodule of $M^{*}$ generated by the columns of $[M^{*}]A^{\text{tr}}$. Because $[M^{*}]A^{\text{tr}}=(A[M])^{\text{tr}}$, we have $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{e}/M_{e})=\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{\prime})$.
Let $U:=X\setminus\mathbb{V}(\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{*}))$. Then the sheaf induced by $M^{*}$ on $U$ is locally free of rank $e$. Because $X$ is reduced so is $U$. If $A$ is generic, by Kleiman’s transversality theorem [Kl74, Rmks. 6 and 7], the analytic subspace in $U$ defined by the ideal sheaf $\mathrm{Fitt}_{0}(M^{*}/M^{\prime})$ is either empty or reduced and of dimension $1$.
Suppose $S^{\prime}$ is nonempty. Let $x\neq 0$ be a point in $S^{\prime}$. Then $x\in\mathbb{V}(\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{\prime}))$ and $x\not\in\mathbb{V}(\mathrm{Fitt}_{l-e}(\mathcal{O}_{X}^{l}/M^{*}))$. Consider the following short exact sequence
(6) $0\longrightarrow M^{*}_{x}/M^{\prime}_{x}\longrightarrow\mathcal{O}_{X,x}^{l}/M^{\prime}_{x}\longrightarrow\mathcal{O}_{X,x}^{l}/M^{*}_{x}\longrightarrow 0$
Because $\mathcal{O}_{X,x}^{l}/M^{*}_{x}$ is free, (6) is a split exact sequence. By [Stks, Tag 07ZA] we have
$\mathrm{Fitt}_{l-e}(\mathcal{O}_{X,x}^{l}/M^{\prime}_{x})=\sum_{i+j=l-e}\mathrm{Fitt}_{i}(M^{*}_{x}/M^{\prime}_{x})\mathrm{Fitt}_{j}(\mathcal{O}_{X,x}^{l}/M^{*}_{x})=\mathrm{Fitt}_{0}(M^{*}_{x}/M^{\prime}_{x})$
because $\mathcal{O}_{X,x}^{l}/M^{*}_{x}$ is free of rank $l-e$ so the only nonzero term in the sum above is the one corresponding to $(i,j)=(0,l-e)$. Thus locally at $x$, the subspace $S^{\prime}$ is defined in $\mathcal{O}_{X,x}$ by $\mathrm{Fitt}_{0}(\mathcal{O}_{X,x}^{e}/(M_{e})_{x})=\mathrm{Fitt}_{l-e}(\mathcal{O}_{X,x}^{l}/M^{\prime}_{x})=\mathrm{Fitt}_{0}(M^{*}_{x}/M^{\prime}_{x})$. So $S^{\prime}$ is reduced at $x$. ∎
Proof of Theorem 1.1
Choose $y$ close enough to 00. Select a generic $A\in\mathrm{Mat}((n-1)\times p,\mathbb{C})$. Denote by $S^{\prime}$ the subscheme of $X$ cut out by $\mathrm{Fitt}_{0}(\mathcal{O}_{X}^{n-1}/J_{\mathrm{rel}}(Z))$. Since $[J_{\mathrm{rel}}(Z)]$ is the relative Jacobian matrix of $Z$, then $S^{\prime}$ is the union along $Y$ of the singular points of $Z_{y}$ in $X_{y}$. As $X_{y}$ is smooth for $y\neq 0$, then $S’_{y}$ is set-theoretically $S_{y}$ the intersection of $W_{y}$ and $X_{y}$.
If $S^{\prime}_{y}$ is empty for $y\neq 0$, then so is $S_{y}$. Thus, by Proposition 3.2 $S$ is empty. So each $X_{y}$ is therefore a complete intersection and $I(X_{y},W_{y})=0$ for all $y$. If $S^{\prime}_{y}$ is nonempty, then so is $S_{y}$. By Prp. 4.2 applied with $M=J_{\mathrm{rel}}(X)$ and $M_{e}=J_{\mathrm{rel}}(Z)$, we get that $S^{\prime}$ is generically reduced. Thus by generic smoothness $S^{\prime}_{y}$ is reduced for all $y\neq 0$ after possibly replacing $Y$ by a smaller neighborhood of $0\in Y$. By Prp. 4.1 $I_{s_{y}}(X_{y},W_{y})=1$ for each $s_{y}\in S_{y}$. Thus $S_{y}$ is reduced, too. So, $S^{\prime}_{y}=S_{y}$ for $y\neq 0$. By Proposition 3.2 and Proposition 4.1 for $y\neq 0$ we have
(7) $I_{0}(X_{0},W_{0})=I(X_{y},W_{y})=\sum_{s_{y}\in S_{y}}I_{s_{y}}(X_{y},W_{y})=\#(S_{y}).$
By Proposition 2.2 $C_{\mathrm{rel}}(X)=\mathrm{Projan}(\mathcal{R}(J_{\mathrm{rel}}(Z))$. Let $B\in\mathrm{Mat}(n\times e,\mathbb{C})$. Consider the $\mathcal{O}_{X}$-module $J_{\mathrm{rel}}(Z)^{\prime}$ generated by the columns of $[J_{\mathrm{rel}}(Z)]B$. Denote the images of $J_{\mathrm{rel}}(Z)^{\prime}$ and $J_{\mathrm{rel}}(Z)$ in $\mathcal{O}_{X_{0}}^{n-1}$ by $J(Z_{0})^{\prime}$ and $J(Z_{0})$, respectively. Choose $B$ generic so that $J(Z_{0})^{\prime}$ is a reduction $J(Z_{0})$ in $\mathcal{O}_{X_{0}}^{n-1}$. Because $J_{\mathrm{rel}}(Z)^{\prime}$ is generated by $n-1$ elements, each of which defines a generic hyperplane ˇ$\check{\mathbb{P}}^{n-1}$, then the ideal $J_{\mathrm{rel}}(Z)^{\prime}\mathcal{R}(J_{\mathrm{rel}}(Z))$ cuts out $\Gamma_{\mathrm{rel}}^{1}(X)$ before the projection to $X$. Thus $\mathrm{Supp}_{X}(J_{\mathrm{rel}}(Z)/J_{\mathrm{rel}}(Z)^{\prime})=\Gamma_{\mathrm{rel}}^{1}(X)$. The genericity condition on $B$ and Kleiman’s transversality theorem ensure that $\Gamma_{\mathrm{rel}}^{1}(X)$ meets $S$ only at 00 after possibly replacing the representative for $(X,0)$ by a smaller one.
Choosing $B$ generic enough by [Kl74, Rmks. 6 and 7] it follows that the principle ideal $(g):=\mathrm{Fitt}_{0}(\mathcal{O}_{X,0}^{n-1}/J_{\mathrm{rel}}(Z)^{\prime})$ defines a reduced subvariety of $X$, which is therefore Cohen-Macaulay, and thus it is flat over $Y$ because $Y$ is smooth. For $y\in Y$ denote the image of $g$ in $\mathcal{O}_{X_{y}}$ by $g_{y}$. By flatness $\dim_{\mathbb{C}}\mathcal{O}_{X_{0}}/(g_{0})=\dim_{\mathbb{C}}\mathcal{O}_{X_{y}}/(g_{y}).$ But $\dim_{\mathbb{C}}\mathcal{O}_{X_{0}}/(g_{0})=e(J(Z_{0}),\mathcal{O}_{X_{0}}^{n-1})$. By Proposition 2.3 we have $e(J(Z_{0}),\mathcal{O}_{X_{0}}^{n-1})=e(\mathrm{Jac}(X_{0}))$. Thus
(8) $e(\mathrm{Jac}(X_{0}))=\dim_{\mathbb{C}}\mathcal{O}_{X_{y}}/(g_{y}).$
It remains to interpret the right-hand side of (8). We need to count the points cut out from $X_{y}$ by $\mathrm{Fitt}_{0}(\mathcal{O}_{X_{y}}^{n-1}/J(Z_{y})^{\prime})$ for $y\neq 0$. There are two disjoint set of points: those that constitute the support of $J_{\mathrm{rel}}(Z_{y})/J_{\mathrm{rel}}(Z_{y})^{\prime}$ and those that are not in it. The first set of points is precisely $\Gamma_{\mathrm{rel}}^{1}(X)_{y}$. By definition the latter set is $S^{\prime}_{y}$ which is the same as $S_{y}$ as determined above. Thus
(9) $\dim_{\mathbb{C}}\mathcal{O}_{X_{y}}/(g_{y})=\#(\Gamma_{\mathrm{rel}}^{1}(X)_{y})+\#(S_{y}).$
Combining (7), (8), (9) and (3) we obtain $e(\mathrm{Jac}(X_{0}))-I_{0}(X_{0},W_{0})=\#(\Gamma_{\mathrm{rel}}^{1}(X)_{y})=\mu+m-1.$
Remark. A consequence of the formula we just proved is that the complete intersection discrepancy $I_{0}(X_{0},W_{0})$ of $X_{0}$ is an intrinsic invariant for a reduced smoothable $(X_{0},0)$.
The approach of selecting a generic complete intersection $Z_{0}$ to define an invariant of $(X_{0},0)$ does not work in dimension greater than $1$, because the components of $Z_{0}$ must intersect in codimension $1$, so $J(Z_{0})$ will not have finite colength in $\mathcal{O}_{X_{0}}^{e}$ where $e=\operatorname{codim}(X_{0},\mathbb{C}^{n}).$
Proof of Corollary 1.2
By Theorem 1.1 we have $e(\mathrm{Jac}(X_{y}))-I_{0}(X_{y},W_{y})=\mu(X_{y})+m(X_{y})-1$, where $\mu(X_{y})$ and $m(X_{y})$ are the Milnor number and the multiplicity of $X_{y}$ at $0\in\mathbb{C}^{n}$. Because $\mu(X_{y})$ and $m(X_{y})$ are upper semi-continuous invariants the function $y\rightarrow e(\mathrm{Jac}(X_{y}))-I_{0}(X_{y},W_{y})$ is constant along $Y$ if and only if $\mu(X_{y})$ and $m(X_{y})$ are constant along $Y$. Note that if $m(X_{y})$ is constant along $Y$, the singular locus does not split, i.e. the union of singularities of the fibers $X_{y}$ is $Y$. The rest of the proof follows from [BGG80, Thm. III.3].
5. Examples
Three lines in $\mathbb{C}^{3}$
Let $X_{0}$ be the union of the three axis of $\mathbb{C}^{3}$ given by the equations $f_{1}=xy,f_{2}=yz,f_{3}=xz.$
The Jacobian matrix of $X_{0}$ is
$\begin{pmatrix}y&x&0\\ 0&z&y\\ z&0&x\end{pmatrix}.$
The Jacobian ideal $\mathrm{Jac}(X_{0})$ in $\mathcal{O}_{X_{0}}$ is generated by $x^{2},y^{2}$ and $z^{2}$. A reduction of it is generated by $x^{2}+y^{2}+z^{2}$. Thus $e(\mathrm{Jac}(X_{0}))=\dim_{\mathbb{C}}\mathbb{C}[x,y,z]/(x^{2}+y^{2}+z^{2},xy,yz,xz)=6$. The multiplicity of $(X_{0},0)$ is $3$. To compute the intersection multiplicity term we chose $Z_{0}=\mathbb{V}(f_{1}+f_{2},f_{1}+f_{3})$ so that $W_{0}=\mathbb{V}(x-y,y+z)$ and thus $I_{0}(X_{0},W_{0})=2$. Thus
$\mu(X_{0})=e(\mathrm{Jac}(X_{0}))-I_{0}(X_{0},W_{0})-m+1=6-2-3+1=2.$
This agrees with [BG80, Lemma 1.2.4]. A smoothing of $X_{0}$ is given by the equations $F_{1}=xy+tx+t^{2},F_{2}=yz+ty+t^{2},F_{3}=zx+tz+t^{2}$, where $X_{t}$ for $t\neq 0$ is a smooth irreducible complete intersection curve. For this family $W$ is the plane $\mathbb{V}(y+z,x+z+2t)$. The line $W_{t}$ for $t\neq 0$ intersects $X_{t}$ in two distinct points.
The $(3,4,5)$ curve in $\mathbb{C}^{3}$
Next we consider the space curve $X_{0}$ given by the following equations $f_{1}=x^{2}y-z^{2},f_{2}=x^{3}-yz,f_{3}=y^{2}-xz$, which come from the $2\times 2$ minors of the matrix $\begin{pmatrix}x&y&z\\ y&z&x^{2}\end{pmatrix}.$ The parametrization for $X_{0}$ is given by $\{(t^{3},t^{4},t^{5})|t\in\mathbb{C}\}\subset\mathbb{C}^{3}$. It is a well known example of a non-planar curve, which is Cohen-Macaulay of codimension $2$, but not Gorenstein. Its Jacobian matrix is
$\begin{pmatrix}2xy&x^{2}&-2z\\ 3x^{2}&-z&-y\\ -z&2y&-x\end{pmatrix}.$
Let $f_{0}=2y^{2}+xz+3yz+z^{2}-4x^{3}-7x^{2}y-2xy^{2}-7x^{2}z-2xyz-3x^{4}$ be a generic combination of the $2\times 2$ minors of the Jacobian matrix. Then (via a Singular computation [DGPS]) $e(\mathrm{Jac}(X_{0}))=\dim_{\mathbb{C}}\mathcal{O}_{X_{0},0}/(f_{0})=8.$
To compute the intersection term use the complete intersection $Z_{0}=\mathbb{V}(f_{1}+f_{2},f_{1}+f_{3})$, which gives a generic complete intersection, such that $e(J(X_{0})_{2},\mathcal{O}_{X_{0}}^{2})=e(\mathrm{Jac}(X_{0}))$, where $\mathcal{O}_{X_{0}}^{2}/J(X_{0})_{2}$ has presentation matrix $[J(X_{0})_{2}]:=\begin{pmatrix}1&1&0\\ 1&0&1\end{pmatrix}[J(X_{0})]$. By a quotient ideal computation we find
$W_{0}=\mathbb{V}(x+y+z,y+z+x^{2})\ \text{and}\ I_{0}(X_{0},W_{0})=\dim_{\mathbb{C}}\mathcal{O}_{\mathbb{C}^{3},0}/I(X_{0})+I(W_{0})=2.$
Thus
$\mu(X_{0})=e(\mathrm{Jac}(X_{0}))-I_{0}(X_{0},W_{0})-m+1=8-2-3+1=4.$
Now let’s choose $Z^{\prime}_{0}=\mathbb{V}(f_{1},f_{3})$. Denote by $J(X_{0})_{2}^{\prime}$ the submodule in $\mathcal{O}_{X_{0}}^{2}$ generated by the columns of the matrix obtained from the Jacobian matrix of $X_{0}$ by erasing the second row. Then
$e(J(X_{0})_{2}^{\prime},\mathcal{O}_{X_{0}}^{2})=9>e(J(X_{0})_{2},\mathcal{O}_{X_{0}}^{2}).$
But $W_{0}^{\prime}=(y,z)$. So $I_{0}(X_{0},W_{0}^{\prime})=3$. A smoothing of $X_{0}$ is given by the $2\times 2$ minors of $\begin{pmatrix}x&y&z\\ y&z&x^{2}+t\end{pmatrix}.$ Then
$X\times_{\mathbb{C}^{4}}W^{\prime}=\mathbb{V}(x,y,z)\cup\mathbb{V}(y,z,x^{2}+t)$
and $X_{t}$ and $W^{\prime}_{t}$ meet transversally at $3$ points for $t\neq 0$. As predicted by Gaffney’s Multiplicity-Polar Theorem (see Theorem 6.1)
$e(J(X_{0})_{2},\mathcal{O}_{X_{0}}^{2})-I_{0}(X_{0},W_{0})=e(J(X_{0})_{2}^{\prime},\mathcal{O}_{X_{0}}^{2})-I_{0}(X_{0},W_{0}^{\prime})=6.$
The BR-multiplicities $e(J(X_{0})_{2},\mathcal{O}_{X_{0}}^{2})$ and $e(J(X_{0})_{2}^{\prime},\mathcal{O}_{X_{0}}^{2})$ are computed through a sum of intersection numbers of the exceptional divisors of the blowup of $\mathrm{Projan}(\mathrm{Sym}(\mathcal{O}_{X_{0}}^{2}))=X_{0}\times\mathbb{P}^{1}$ with respect to the ideals $J(X_{0})_{2}\mathrm{Sym}(\mathcal{O}_{X_{0}}^{2})$ and $J(X_{0})_{2}^{\prime}\mathrm{Sym}(\mathcal{O}_{X_{0}}^{2})$ ([KT94, Sct. 5]). The reason $e(J(X_{0})_{2},\mathcal{O}_{X_{0}}^{2})\neq e(J(X_{0})_{2}^{\prime},\mathcal{O}_{X_{0}}^{2})$ is that although $\mathrm{Projan}(\mathcal{R}(J(X_{0})_{2}))=\mathrm{Projan}(\mathcal{R}(J(X_{0})_{2}^{\prime}))$ the two blowups are not isomorphic.
Finally, observe that for the special choice $Z^{\prime\prime}_{0}=\mathbb{V}(f_{1},f_{2})$, in the induced deformation of $Z^{\prime\prime}$ by the smoothing of $X_{0}$ above, we have $W^{\prime\prime}=\mathbb{V}(z,x^{2}-t)$ and $X\times_{\mathbb{C}^{4}}W^{\prime\prime}=\mathbb{V}(z,y^{2},x^{2}-t)$. For $t\neq 0$, one finds that $W_{t}^{\prime\prime}$ consists of two lines each tangent to $X_{t}$, i.e. $X_{t}$ and $W^{\prime\prime}_{t}$ do not intersect transversally.
A Whitney equisingular family
The following example is taken from [BG80, Sct. 7]. Consider the one-parameter family of irreducible curves $(X,0)\rightarrow(D,0)$ given parameterically by $\mathbb{C}\{u^{4},u^{7}+tu^{6},u^{9},u^{10}\}$, where $t$ is the uniformazing parameter of $\mathcal{O}_{D}$. Using Singular we compute that the base space of miniversal deformations is irreducible and that $X_{0}$ is smoothable. Thus $X_{t}$ is smoothable for each $t$. We compute that for each $t$ the delta invariant of $X_{t}$ is $5$ and so $\mu(X_{t},0)=10$ (in [BG80] it is wrongly claimed that $\mu(X_{t},0)=12$). Thus, for each $t\in D$ we have
$e(\mathrm{Jac}(X_{t},0))-I_{0}(X_{t},W_{t})=\mu(X_{t},0)+m(X_{t},0)-1=10+4-1=13.$
However, a computation with Singular reveals that $e(\mathrm{Jac}(X_{0}))=21$ and $I_{0}(X_{0},W_{0})=8$, whereas $e(\mathrm{Jac}(X_{t},0))=19$ and $I_{0}(X_{t},W_{t})=6$ for $t\neq 0$. This computation shows that apart from the case of families of complete intersection curves, $e(\mathrm{Jac}(X_{t},0))$ is not a Whitney equisingularity invariant in general. So the presence of the correcting term $I_{0}(X_{t},W_{t})$ to $e(\mathrm{Jac}(X_{t},0))$ is necessary for obtaining numerical control of Whitney equisingularity.
6. Appendix
In the proof of Theorem 1.1 we implicitly used a special case of Gaffney’s Multiplicity-Polar Theorem. Here we give a new proof of this particular version of his result inspired by our proof of Theorem 1.1.
Let $h\colon(X,x_{0})\rightarrow(Y,y_{0})$ be a morphism with equidimensional fibers of positive dimension $d$ between equidimensional complex analytic varieties such that $X$ is generically reduced and $Y$ is smooth of dimension $k$. Let $N\subset F$ be coherent $\mathcal{O}_{X}$-modules such that $F$ is free of rank $e$. Suppose $T:=\mathrm{Supp}_{X}(F/N)$ is finite over $Y$. Let $\mathcal{R}(N)$ be the Rees algebra of $N$ it is the subalgebra of $\mathrm{Sym}(F)$ generated in degree one by the generators of $N$.
For each $y\in Y$ denote by $F(y)$ the restriction of $F$ to $X_{y}$ and by $N(y)$ the image of $N$ in $F(y)$. As before, define the Buchsbaum–Rim multiplicity $e(N(y),F(y))$ as the normalized leading coefficient of $\dim_{\mathbb{C}}(F^{l}(y)/N^{l}(y))$ which is a polynomial of degree $r:=d+e-1$ for $l$ large enough, where $F^{l}(y)$ and $N^{l}(y)$ are the $l$th graded components of $\mathrm{Sym}(F(y))$ and the Rees algebra $\mathcal{R}(N(y))$, respectively. Note that $\mathrm{Projan}(\mathcal{R}(N))\subset X\times\mathbb{P}^{g(N)-1}$ where $g(N)$ is the cardinality of a generating set for $N$ as an $\mathcal{O}_{X}$-module. Denote by $\pi\colon\mathrm{Projan}(\mathcal{R}(N))\rightarrow X$ the structure morphism. Consider the composition of maps
$\pi^{-1}(T)\hookrightarrow X\times\mathbb{P}^{g(N)-1}\xrightarrow{pr_{2}}\mathbb{P}^{g(N)-1}.$
As $N$ is generically free of rank $e$, by Kleiman’s Transversality Theorem [Kl74], the intersection of $\pi^{-1}(T)$ with a general plane $H_{r}$ from $\mathbb{P}^{g(N)-1}$ of codimension $r$ is of dimension at most $\dim Y-1$. Denote by $\Gamma^{k}(N)$ the projection of $\mathrm{Proj}(\mathcal{R}(N))\cap H_{r}$ to $X$. This is what Gaffney [Gaff04] calls the $k$-dimensional polar variety of $N$. For $y$ in a Zariski open subset $U$ of $Y$, the fiber of $\Gamma^{k}(N)$ over $y$ consists of the same number of points, each of them appearing with multiplicity one because $X$ is generically reduced, and because locally at each one of them $N$ is free. Denote this number by $\mathrm{deg}_{Y}\Gamma^{k}(N)$. The following is a special case of Gaffney’s Multiplicity-Polar Theorem (see [Gaf07] and [Ran22, Sct. 2] for generalizations).
Theorem 6.1 (Gaffney).
Suppose $X$ is Cohen–Macaulay and $T$ is finite over $Y$. Then for each $y$ in Zariski open subset $U$ in $Y$ we have
$e(N(y_{0}),F(y_{0}))-e(N(y),F(y))=\mathrm{deg}_{Y}\Gamma^{k}(N).$
Proof.
We follow the proof of Theorem 1.1. Let $N^{\prime}$ be a submodule of $N$ generated by $r$ generic linear combinations of generators for $N$ such that $N^{\prime}(y_{0})$ is a reduction of $N(y_{0})$. Set $\mathcal{I}:=\mathrm{Fitt}_{0}(F/N^{\prime})$. Denote by $T^{\prime}$ the subspace of $X$ defined by $\mathcal{I}$. Because $\mathrm{Supp}(F(y_{0})/N^{\prime}(y_{0}))$ is finite and because $T\subset\mathrm{Supp}(F/N^{\prime})$, then $\dim T^{\prime}=k$. Since the codimension of $\mathcal{O}_{X}/\mathcal{I}$ is right, then $T^{\prime}$ is determinantal, and so it is Cohen–Macaulay because $X$ is. Because $Y$ is smooth, then $T^{\prime}\rightarrow Y$ is flat. Thus $\dim_{\mathbb{C}}(\mathcal{O}_{X_{y_{0}}}/\mathcal{I}(y_{0}))=\dim_{\mathbb{C}}(\mathcal{O}_{X_{y}}/\mathcal{I}(y))$ for each $y\in Y$. Also, $h\colon(X,x_{0})\rightarrow(Y,y_{0})$ is flat, because $X$ is Cohen–Macaulay, $X$ and the fibers of $h$ are equidimensional, and $Y$ is smooth. Thus $X_{y}$ is Cohen-Macaulay for each $y\in Y$. By [BR64, 4.5, p. 223] $e(N(y_{0}),F(y_{0}))=\dim_{\mathbb{C}}(\mathcal{O}_{X_{y_{0}}}/\mathcal{I}(y_{0}))$. It remains to interpret $\dim_{\mathbb{C}}(\mathcal{O}_{X_{y}}/\mathcal{I}(y))$ for generic $y\in Y$.
Set $T^{\prime\prime}:=\mathrm{Supp}(N/N^{\prime})$. Note that set-theoretically, $T^{\prime}=T\cup T^{\prime\prime}$. The generators of $N^{\prime}$ give the ideal of the plane $H_{r}$ in $\mathbb{P}^{g(N)-1}$ used to define $\Gamma^{k}(N)$. Denote by $U$ the complement in $Y$ of the Zariski closure of $h\circ\pi(\pi^{-1}(T)\cap H_{r})$. Then $T^{\prime\prime}_{y}=\Gamma^{k}(N)_{y}$ for $y\in U$. By construction $T_{y}$ and $T^{\prime\prime}_{y}$ are disjoint for $y\in U$. Thus
$\dim_{\mathbb{C}}(\mathcal{O}_{X_{y}}/\mathcal{I}(y))=\sum_{t_{y}\in T_{y}}\dim_{\mathbb{C}}(\mathcal{O}_{X_{y},t_{y}}/\mathcal{I}(y,t_{y}))+\mathrm{deg}_{Y}\Gamma^{k}(N)$
where $\mathcal{I}(y,t_{y})$ is the image of $\mathcal{I}$ in $\mathcal{O}_{X_{y},t_{y}}$. Because the formation of Fitting ideals commutes with base change and because $\mathcal{O}_{X_{y},t_{y}}$ is Cohen-Macaulay, [BR64, 4.5, p. 223] gives $e(N(y),F(y))=\sum_{t_{y}\in T_{y}}\dim_{\mathbb{C}}(\mathcal{O}_{X_{y},t_{y}}/\mathcal{I}(y,t_{y}))$. The proof of the theorem is now complete. ∎